Focus groups of 14 people are randomly selected to discuss products of the Yummy Company. It is determined that the mean number (per group) who recognize the Yummy brand name is 8.9, and the standard deviation is 0.69. Would it be unusual to randomly select 14 people and find that greater than 12 recognize the Yummy brand name?
Answer by stanbon(75887)

Focus groups of 14 people are randomly selected to discuss products of the Yummy Company.
It is determined that the mean number (per group) who recognize the Yummy brand name is 8.9, and the standard deviation is 0.69.
Would it be unusual to randomly select 14 people and find that greater than 12 recognize the Yummy brand name?
------------------------------------------
z(12) = (12-14)/0.69 = -2.8986
---
P(xbar > 12) = P(z > -2.8986) = 0.9981
---
It would not be at all unusual to find that greater than 12
recognize the brand name.
==================================
Cheers,
Stan H.